# Non-Iterative Coordination of Interconnected Power Grids via Dimension-Decomposition-Based Flexibility Aggregation

### v2

@@ III-B Generation Units Constraints
The operational constraints for all renewable generation station $i\in\mathcal{I}^{\text{rgu}}$ are as follows
@@
where (11a) and (11b) are the power and ramp rate constraints, respectively. (11c) and (11d) are the reserve constraints. $Rd^{\text{req}}_{t}$ and $Ru^{\text{req}}_{t}$ denote the total down and up reserve requirements of the entire RPG at time $t$, as determined by the dispatch center based on the fluctuation in the generation of renewable energy and load power [25].
@@ III-C N-1 Security Constraints
where the subscript $c$ represents the index of $N$-1 contingency scenarios, and $\mathcal{C}$ is the index set of key contingency. $\Delta\bm{p}^{\text{tpg}}$, $\Delta\bm{p}^{\text{rgu}}$ and $\Delta\bm{\theta}$ denote the maximum allowable variations of the power output of thermal power generators and renewable generation stations, as well as the voltage angles of buses in the contingency scenarios, respectively.
@@ III-E Model Integration
The internal variables $\bm{x}^{\text{int}}$ include all the variables of power generation units, and network states, under both normal and $N$-1 contingency operation conditions.
@@
This choice of boundary variables is based on the fact that the transmission power of the DC tie-lines is directly controlled by converters, whereas the AC tie-line power transmission depends on the voltage angles of the boundary buses [4], governed by power balance and network constraints
@@
where $\theta_{r,i,t}^{\text{bd}}$ and $\theta_{s,j,t}^{\text{bd}}$ denote the voltage angles of the $i$-th bus in region $r$ and the $j$-th bus in region $s$ at time $t$, respectively. $p_{r,ij,t}^{\text{tie}}$ and $p_{s,ji,t}^{\text{tie}}$ represent power flows from region $r$ to region $s$ through tie-line $ij$ and vice versa, respectively, and $x_{ij}$ is the reactance of this AC tie-line $ij$.
@@ IV-A Dimension-Decomposition-Based Calculation of Aggregated Flexibility Region
This section presents the proposed dimension-decomposition-based method for calculating the aggregated flexibility region of an RPG. The process consists of five key steps: 1) Analyze the temporal and spatial coupling relationships of boundary variables; 2) Decompose the flexibility region into a series of lower-dimensional subspaces using dimensionality reduction projection; 3) Approximate the projected regions within these subspaces using different shape templates; 4) Reconstruct the overall inner-approximated flexibility region in $\operatorname{space}(\bm{x}^{\text{bd}})$ by recombining the inner-approximated regions in the subspaces; and 5) Adjust the parameters to ensure that the inner approximation condition ^$\hat{\mathcal{R}}_{\text{L}}\subseteq\mathcal{R}_{\text{L}}$ is satisfied. A detailed explanation of each step follows, with an illustrative diagram provided in Fig. 6.
@@ IV-A1 Analysis of Boundary Variables
To calculate the aggregated flexibility region $\mathcal{R}_{\text{L}}$ of an RPG in $\operatorname{space}(\bm{x}^{\text{bd}})$, it is crucial to understand the coupling relationships between the boundary variables $\bm{x}^{\text{bd}}$ first. By analyzing the coupling constraints of the boundary variables in (8)-(12), all the constraints can be categorized into two types. Accordingly, all the boundary variables can be categorized into two groups of variable sets, as illustrated in Fig. 5.
Figure 5: The coupling constraints of boundary variables for a regional power grid.
The first category, temporal coupling constraints, addresses the active output power constraints across all time slots for a single tie-line, denoted as
@@
This form of coupling is mainly driven by generator operation constraints (11) and power balance constraints (9). Consequently, the active output power across time slots for a single tie-line exhibits temporal coupling characteristics, similar to those observed in thermal power generators, which are regulated by both power and ramp rate constraints.
The second category, single time slice coupling constraints, encompasses coupling constraints for boundary variables within each individual time slice $t$, denoted as
@@
These relationships are governed by constraints that apply independently within each time slice, such as network constants (8b), power balance (9), and $N$-1 security constraints (12).
@@
IV-A2 Dimension Decomposition
The coupling relationships of the boundary variables in an RPG can be divided into two categories. Accordingly, the aggregated flexibility region $\mathcal{R}_{\text{L}}\subset\operatorname{space}(\bm{x}^{\text{bd}})$ can, from a geometric perspective, be decomposed into two corresponding groups of subspaces: $\operatorname{space}(\bm{x}^{\text{bd}}_{t})$ for $t\in\mathcal{T}$, and $\operatorname{space}(\bm{p}^{\text{tie}}_{i})$ for $i\in\mathcal{I}^{\text{tie}}$.
To express the coupling relationships within these subspaces, the aggregated flexibility region $\mathcal{R}_{\text{L}}$ can be projected onto lower-dimensional subspaces. Denote the projected polytopes as $\mathcal{E}^{\text{bd}}_{t}\subset\operatorname{space}(\bm{x}^{\text{bd}}_{t})$ for $t\in\mathcal{T}$, and $\mathcal{G}^{\text{tie}}_{i}\subset\operatorname{space}(\bm{p}^{\text{tie}}_{i})$ for $i\in\mathcal{I}^{\text{tie}}$, respectively. An illustrated diagram is shown in Fig. 6(a).
IV-A3 Inner Approximation of Projected Polytopes in the Subspaces
To explicitly calculate the expression of coupling constraints in each subspace, the polytopes $\mathcal{E}^{\text{bd}}_{t}$ and $\mathcal{G}^{\text{tie}}_{i}$ are inner-approximated, respectively, as shown in Fig. 6(b).
@@ IV-A3 Inner Approximation of Projected Polytopes in the Subspaces
where $\overline{\bm{p}}^{\text{tie}}_{i}:=\operatorname{col}_{t}(\overline{p}^{\text{tie}}_{i,t})$, $\underline{\bm{p}}^{\text{tie}}_{i}:=\operatorname{col}_{t}(\underline{p}^{\text{tie}}_{i,t})$, $\overline{\bm{r}}^{\text{tie}}_{i}:=\operatorname{col}_{t}(\overline{r}^{\text{tie}}_{i,t})$, $\underline{\bm{r}}^{\text{tie}}_{i}:=\operatorname{col}_{t}(\underline{r}^{\text{tie}}_{i,t})$. $\bm{b}^{\text{tie}}_{i}$ can be calculated based on the bound shrinking algorithm [16] introduced in Section II-B. Then, (18) can be rewritten as the following compact matrix form:
@@
The single time slice coupling constraints are governed by the complex coupling relationships among the coupling variables that lack an explicit expression, the polytopes $\mathcal{E}^{\text{bd}}_{t}$ can be inner-approximated by the ellipsoidal shape templates, denoted as ^$\hat{\mathcal{E}}^{\text{bd}}_{t}$.
@@
The parameters $\bm{E}_{t}^{\text{bd}}$ and $\bm{e}_{t}^{\text{bd}}$ can be calculated based on the methods introduced in Section II-B.
Notably, the independence of projection and inner approximation calculations in each subspace allows the computation of parameters—$\bm{b}^{\text{tie}}_{i}$ for $i\in\mathcal{I}^{\text{tie}}$, as well as $\bm{E}_{t}^{\text{bd}}$ and $\bm{e}_{t}^{\text{bd}}$ for $t\in\mathcal{T}$—to be efficiently accelerated through parallel computing.
IV-A4 Dimension Recombination
@@
where the constant matrices $\bm{S}^{\text{tie}}_{i}$ and $\bm{H}^{\text{bd}}_{t}$ are used to select the elements $\bm{p}^{\text{tie}}_{i}$ and $\bm{x}^{\text{bd}}_{t}$ from all the boundary variables $\bm{x}^{\text{bd}}$ as defined in (16) and (17), respectively. The recombination region is an approximation of $\mathcal{R}_{\text{L}}$, denoted as ~$\tilde{\mathcal{R}}_{\text{L}}$, where the parameters $\tilde{\bm{b}}^{\text{tie}}_{i}$ and $\tilde{\bm{E}}_{t}^{\text{bd}}$ still need further adjustments in the following step.
IV-A5 Parameters Adjustments
To finally obtain the inner-approximated ^$\hat{\mathcal{R}}_{\text{L}}\subseteq\mathcal{R}_{\text{L}}$, further adjustments are required to the parameters of ~$\tilde{\mathcal{R}}_{\text{L}}$. The parameters adjustment can be achieved iteratively by shrinking the bounds of $\hat{\mathcal{G}}^{\text{tie}}_{i}(\tilde{\bm{b}}^{\text{tie}}_{i})$ and ^$\hat{\mathcal{E}}_{t}^{\text{bd}}(\tilde{\bm{E}}_{t}^{\text{bd}},\bm{e}_{t}^{\text{bd}})$ that make up ~$\tilde{\mathcal{R}}_{\text{L}}$. In each iteration, the parameters are updated to shrink the flexibility region. Denote the parameters in the $k$-th iteration as $\bm{b}^{\text{tie}}_{i(k)}$ and $\bm{E}_{t(k)}^{\text{bd}}$, and the flexibility region in the $k$-th iteration as ^$\hat{\mathcal{R}}_{\text{L}(k)}$. The iteration stops until the condition ^$\hat{\mathcal{R}}_{\text{L}(k)}\subseteq\mathcal{R}_{\text{L}}$ is met. The initial values of the parameters in the iteration can be obtained in the previous dimension recombination step, denoted as $\bm{b}^{\text{tie}}_{i(0)}:=\tilde{\bm{b}}^{\text{tie}}_{i}$ and $\bm{E}_{t(0)}^{\text{bd}}:=\tilde{\bm{E}}_{t}^{\text{bd}}$.
@@ IV-A5 Parameters Adjustments
In this Stackelberg game, the leader variables, $\bm{x}^{\text{bd}}$, aims to guarantee the feasibility of the problem, while the follower variables, $\bm{x}^{\text{int}}$, seeks to make it infeasible. If the follower’s maximization problem is infeasible, it indicates the existence of an outlier $\bm{x}^{\text{out}}_{(k)}$, such that $\bm{x}^{\text{out}}_{(k)}\in\hat{\mathcal{R}}_{\text{L}(k)}$ but $\bm{x}^{\text{out}}_{(k)}\notin\mathcal{R}_{\text{L}}$. In this case, the optimal objective value of (23) is $\infty$- ∞. Otherwise, if the follower’s maximization problem is feasible, no such outlier exists, confirming that ^$\hat{\mathcal{R}}_{\text{L}(k)}\subseteq\mathcal{R}_{\text{L}}$, and the optimal objective value is 0. The problem (23) can be converted into an MILP formulation and solved, as described in detail in the supplemental file [27].
@@
Then update $k\leftarrow k+1$ and proceed to solve (23) to check for the existence of outliers in ^$\hat{\mathcal{R}}_{\text{L}(k+1)}$.
@@
4:  Solve (25) to select the indices of the bounds of ^$\hat{\mathcal{R}}_{\text{L}(k)}$ that $\bm{x}^{\text{out}}_{(k)}$ resides.
@@
Finally, the aggregated flexibility region of RPG is obtained as follows:
^$\hat{\mathcal{R}}_{\text{L}}:=\left\{\bm{x}^{\text{bd}}\left\lvert\,\begin{array}[]{c}\bm{A}^{\text{tie}}_{i}\bm{S}^{\text{tie}}_{i}\bm{x}^{\text{bd}}\leq\bm{b}^{\text{tie}}_{i},\forall i\in\mathcal{I}^{\text{tie}}\\ \left\|\left(\bm{E}_{t}^{\text{bd}}\right)^{-1}\left(\bm{H}^{\text{bd}}_{t}\bm{x}^{\text{bd}}-\bm{e}_{t}^{\text{bd}}\right)\right\|\leq 1,\forall t\in\mathcal{T}\end{array}\right.\right\}$ (28)

### v3

@@ III-B Generation Units Constraints
The operational constraints for all renewable generation station $i\in\mathcal{I}^{\text{rgu}}$ are
@@
where (11a) and (11b) are the power and ramp rate constraints, respectively. (11c) and (11d) are the reserve constraints. $Rd^{\text{tpg}}_{i,t},Ru^{\text{tpg}}_{i,t}$ denote down and up reserve of $i$-th thermal power generator at time $t$. $Rd^{\text{req}}_{t}$ and $Ru^{\text{req}}_{t}$ denote the total down and up reserve requirements of the entire RPG at time $t$, determined by the dispatch center based on the fluctuation in the generation of renewable energy and load power [25].
@@ III-C N-1 Security Constraints
where the subscript $c$ represents the index of $N$-1 contingency scenarios, and $\mathcal{C}$ is the index set of key contingency. $\Delta\bm{p}^{\text{tpg}}$, $\Delta\bm{p}^{\text{rgu}}$ and $\Delta\bm{\theta}$ denote the maximum allowable variations of the power output of thermal power generators and renewable generation stations, as well as the voltage angles of the buses in the contingency scenarios, respectively.
@@ III-E Model Integration
The internal variables $\bm{x}^{\text{int}}$ include all the variables of power generation units and network states, under normal and $N$-1 contingency operation conditions.
@@
This choice of boundary variables is based on the fact that the transmission power of the DC tie-lines is directly controlled by converters, while the AC tie-line power transmission depends on the voltage angles of the boundary buses [4], governed by power balance and network constraints.
@@
where $\theta_{r,i,t}^{\text{bd}}$ and $\theta_{s,j,t}^{\text{bd}}$ denote the voltage angles of the $i$-th bus in region $r$ and the $j$-th bus in region $s$ at time $t$, respectively. $p_{r,ij,t}^{\text{tie}}$ and $p_{s,ji,t}^{\text{tie}}$ represent power flows from region $r$ to region $s$ through the tie-line $ij$ and vice versa, respectively, and $x_{ij}$ is the reactance of this AC tie-line $ij$.
@@ IV-A Dimension-Decomposition-Based Calculation of Aggregated Flexibility Region
This section presents the proposed dimension-decomposition-based method for calculating the aggregated flexibility region of an RPG. There are four main steps in the process. An illustrative diagram of each step is provided in Fig. 7.
To illustrate the method, we use an IEEE 118-bus RPG example, introduced step by step alongside the methodology. This system includes two tie-lines: an AC tie-line connected to Bus 17 and a DC tie-line connected to Bus 31. The schematic diagram of this system is depicted in Fig. 5.
Figure 5: Schematic diagram of the IEEE 118 regional power grid.
@@ IV-A1 Analysis of Boundary Variables
To calculate the aggregated flexibility region $\mathcal{R}_{\text{L}}$ of an RPG in $\operatorname{space}(\bm{x}^{\text{bd}})$, it is crucial to understand the coupling relationships between the boundary variables $\bm{x}^{\text{bd}}$. By analyzing the coupling constraints of the boundary variables in (8)-(12), all the constraints can be categorized into two types. Consequently, all the boundary variables can be categorized into two categories of variable sets, as illustrated in Fig. 6.
Figure 6: The coupling constraints of boundary variables for a regional power grid and two categories of constraints.
The first category, temporal coupling constraints, addresses the active output power constraints across all time slots for a single tie-line. The corresponding variables in each group are denoted as
@@
This form of coupling is mainly driven by generator operational constraints (11) and power balance constraints (9). The active output power across time slots for a single tie-line exhibits temporal coupling characteristics, similar to those observed in thermal power generators, which are regulated by both power and ramp rate constraints.
The second category, single time slice coupling constraints, encompasses coupling constraints for boundary variables within each individual time slice $t$. The corresponding variables in each group are denoted as
@@
They are governed by constraints that apply independently within each time slice, such as network constants (8b), power balance (9), and $N$-1 security constraints (12).
In terms of the IEEE 118 RPG, the boundary variables include the output power of two tie-lines ($p^{\text{tie}}_{17,t}$, $p^{\text{tie}}_{31,t}$) and the voltage angle of Bus 17 ($\theta^{\text{bd}}_{17,t}$) across all time slots $t\in\mathcal{T}$. Accordingly, the first category includes the output power constraints of each tie-line across all time slots. The corresponding two groups of variables are $\bm{p}^{\text{tie}}_{17}:=\operatorname{col}_{t\in\mathcal{T}}(p^{\text{tie}}_{17,t})$ and $\bm{p}^{\text{tie}}_{31}:=\operatorname{col}_{t\in\mathcal{T}}(p^{\text{tie}}_{31,t})$, respectively.
The second category includes the constraints of the output power of two tie-lines and the voltage angle of Bus 17 within each time slot $t\in\mathcal{T}$. The corresponding $T$ groups of variables are $\bm{x}^{\text{bd}}_{t}:=\left[p^{\text{tie}}_{17,t},p^{\text{tie}}_{31,t},\theta^{\text{bd}}_{17,t}\right]^{\top}$.
@@
IV-A2 Dimension Decomposition Projection and Inner-Approximation
Based on the categorization results of all the boundary variables in the previous step, all the coupling relationships of the boundary variables in an RPG can be divided into two categories accordingly. From a geometric perspective, the aggregated flexibility region $\mathcal{R}_{\text{L}}\subset\operatorname{space}(\bm{x}^{\text{bd}})$ can be decomposed into two corresponding groups of subspaces: $\operatorname{space}(\bm{p}^{\text{tie}}_{i})$ for $i\in\mathcal{I}^{\text{tie}}$ and $\operatorname{space}(\bm{x}^{\text{bd}}_{t})$ for $t\in\mathcal{T}$.
To express the coupling relationships within these subspaces, the aggregated flexibility region $\mathcal{R}_{\text{L}}$ can be projected onto lower-dimensional subspaces. Denote the projected polytopes by $\mathcal{G}^{\text{tie}}_{i}\subset\operatorname{space}(\bm{p}^{\text{tie}}_{i})$ for $i\in\mathcal{I}^{\text{tie}}$, and $\mathcal{E}^{\text{bd}}_{t}\subset\operatorname{space}(\bm{x}^{\text{bd}}_{t})$ for $t\in\mathcal{T}$, respectively. An illustrated diagram is shown in Fig. 7(a). Subsequently, to obtain the explicit expression of the projected polytopes $\mathcal{E}^{\text{bd}}_{t}$ and $\mathcal{G}^{\text{tie}}_{i}$, we use the methods provided in Section II-B to inner-approximate them, as demonstrated in Fig. 7(b).
@@ IV-A2 Dimension Decomposition Projection and Inner-Approximation
where $\overline{\bm{p}}^{\text{tie}}_{i}:=\operatorname{col}_{t}(\overline{p}^{\text{tie}}_{i,t})$, $\underline{\bm{p}}^{\text{tie}}_{i}:=\operatorname{col}_{t}(\underline{p}^{\text{tie}}_{i,t})$, $\overline{\bm{r}}^{\text{tie}}_{i}:=\operatorname{col}_{t}(\overline{r}^{\text{tie}}_{i,t})$, $\underline{\bm{r}}^{\text{tie}}_{i}:=\operatorname{col}_{t}(\underline{r}^{\text{tie}}_{i,t})$. $\bm{b}^{\text{tie}}_{i}$ can be calculated based on Section II-B. Then, (18) can be rewritten as the following compact matrix form:
@@
The single time slice coupling constraints are governed by the complex coupling relationships among the coupling variables that lack an explicit expression; the polytopes $\mathcal{E}^{\text{bd}}_{t}$ can be inner-approximated by the ellipsoidal shape templates, denoted as ^$\hat{\mathcal{E}}^{\text{bd}}_{t}$.
@@
The parameters $\bm{E}_{t}^{\text{bd}}$ and $\bm{e}_{t}^{\text{bd}}$ can be calculated based on Section II-B.
In particular, the independence of projection and inner approximation calculations in each subspace allows the computation of parameters, $\bm{b}^{\text{tie}}_{i}$ for $i\in\mathcal{I}^{\text{tie}}$, as well as $\bm{E}_{t}^{\text{bd}}$ and $\bm{e}_{t}^{\text{bd}}$ for $t\in\mathcal{T}$, to be efficiently accelerated by parallel computing.
As for the IEEE 118 RPG case, the temporal coupling flexibility region across all time slots including the power bounds ($\overline{\bm{p}}^{\text{tie}}_{17}$, $\underline{\bm{p}}^{\text{tie}}_{17}$, $\overline{\bm{p}}^{\text{tie}}_{31}$, $\underline{\bm{p}}^{\text{tie}}_{31}$) and ramp rate bounds ($\overline{\bm{r}}^{\text{tie}}_{17}$, $\underline{\bm{r}}^{\text{tie}}_{17}$, $\overline{\bm{r}}^{\text{tie}}_{31}$, $\underline{\bm{r}}^{\text{tie}}_{31}$) of these two tie-lines are shown in Fig. 8.
Figure 8: The aggregated parameters of two tie-lines in the IEEE 118 RPG: (a) Power bounds and (b) Ramp rate bounds.
In addition, taking the time slot $t_{0}=12:00$ as an example, the coupling relationship between the boundary variables $\bm{x}^{\text{bd}}_{t_{0}}:=[p^{\text{tie}}_{17,t_{0}},p^{\text{tie}}_{31,t_{0}},\theta^{\text{bd}}_{17,t_{0}}]^{\top}$ is shown in Fig. 9. The ellipsoid represents the inner-approximated flexibility region ^$\hat{\mathcal{E}}_{t_{0}}^{\text{bd}}$ with an ellipsoidal shape template. The gray polytope represents the actual flexibility region $\mathcal{E}^{\text{bd}}_{t_{0}}$ obtained by searching for all extreme points. In practice, the RPG may have far more tie-lines, making it intractable to find all the extreme points of the actual flexibility region like this case. Therefore, the ellipsoidal shape template is a convenient tool to self-adaptively inner-approximate the flexibility region with fewer parameters.
IV-A3 Dimension Recombination
@@
where the constant matrices $\bm{S}^{\text{tie}}_{i}$ and $\bm{H}^{\text{bd}}_{t}$ are used to select the elements $\bm{p}^{\text{tie}}_{i}$ and $\bm{x}^{\text{bd}}_{t}$ from all the boundary variables $\bm{x}^{\text{bd}}$ defined in (16) and (17), respectively. The recombination region is an approximation of $\mathcal{R}_{\text{L}}$, denoted as ~$\tilde{\mathcal{R}}_{\text{L}}$, where the parameters $\tilde{\bm{b}}^{\text{tie}}_{i}$ and $\tilde{\bm{E}}_{t}^{\text{bd}}$ still need further adjustments in the following step.
IV-A4 Parameter Adjustment
To finally obtain the inner-approximated ^$\hat{\mathcal{R}}_{\text{L}}\subseteq\mathcal{R}_{\text{L}}$, further adjustments are required to the parameters of ~$\tilde{\mathcal{R}}_{\text{L}}$. Parameter adjustment can be achieved iteratively by shrinking the bounds of $\hat{\mathcal{G}}^{\text{tie}}_{i}(\tilde{\bm{b}}^{\text{tie}}_{i})$ and ^$\hat{\mathcal{E}}_{t}^{\text{bd}}(\tilde{\bm{E}}_{t}^{\text{bd}},\bm{e}_{t}^{\text{bd}})$ that make up ~$\tilde{\mathcal{R}}_{\text{L}}$. In each iteration, the parameters are updated to shrink the flexibility region. Denote the parameters in the $k$-th iteration as $\bm{b}^{\text{tie}}_{i(k)}$ and $\bm{E}_{t(k)}^{\text{bd}}$, and the flexibility region in the $k$-th iteration as ^$\hat{\mathcal{R}}_{\text{L}(k)}$. The iteration stops until the condition ^$\hat{\mathcal{R}}_{\text{L}(k)}\subseteq\mathcal{R}_{\text{L}}$ is met. The initial values of the parameters in the iteration can be obtained in the previous dimension recombination step, denoted $\bm{b}^{\text{tie}}_{i(0)}:=\tilde{\bm{b}}^{\text{tie}}_{i}$ and $\bm{E}_{t(0)}^{\text{bd}}:=\tilde{\bm{E}}_{t}^{\text{bd}}$.
@@ IV-A4 Parameter Adjustment
In this Stackelberg game, the leader variables, $\bm{x}^{\text{bd}}$, aim to guarantee the feasibility of the problem, while the follower variables, $\bm{x}^{\text{int}}$, aim to make it infeasible. If the follower’s maximization problem is infeasible, it indicates the existence of an outlier $\bm{x}^{\text{out}}_{(k)}$, such that $\bm{x}^{\text{out}}_{(k)}\in\hat{\mathcal{R}}_{\text{L}(k)}$ but $\bm{x}^{\text{out}}_{(k)}\notin\mathcal{R}_{\text{L}}$. In this case, the optimal objective value of (23) is $\infty$- ∞. Otherwise, if the follower’s maximization problem is feasible, no such outlier exists, confirming that ^$\hat{\mathcal{R}}_{\text{L}(k)}\subseteq\mathcal{R}_{\text{L}}$, and the optimal objective value is 0. The problem (23) can be converted into an MILP formulation and solved, as described in detail in the supplementary file [28].
@@
Then update $k\leftarrow k+1$ and solve (23) again to check for any outliers in ^$\hat{\mathcal{R}}_{\text{L}(k+1)}$.
@@
4:  Solve (25) identify which bounds of ^$\hat{\mathcal{R}}_{\text{L}(k)}$ the outlier $\bm{x}^{\text{out}}_{(k)}$ lies on.
@@
The outputs of this algorithm are the adjusted vector $\bm{b}^{\text{tie}}_{i}$ and matrix $\bm{E}_{t}^{\text{bd}}$ of the inner-approximated flexibility region ^$\hat{\mathcal{R}}_{\text{L}}$. Finally, the aggregated flexibility region of RPG is obtained as follows:
^$\hat{\mathcal{R}}_{\text{L}}:=\left\{\bm{x}^{\text{bd}}\left\lvert\,\begin{array}[]{c}\bm{p}^{\text{tie}}_{i}:=\bm{S}^{\text{tie}}_{i}\bm{x}^{\text{bd}}\in\hat{\mathcal{G}}^{\text{tie}}_{i},\forall i\in\mathcal{I}^{\text{tie}}\\ \bm{x}_{t}^{\text{bd}}:=\bm{H}^{\text{bd}}_{t}\bm{x}^{\text{bd}}\in\hat{\mathcal{E}}_{t}^{\text{bd}},\forall t\in\mathcal{T}\end{array}\right.\right\}$ (28)
where
$\hat{\mathcal{G}}^{\text{tie}}_{i}:=\left\{\bm{p}^{\text{tie}}_{i}\left\lvert\,\bm{A}^{\text{tie}}_{i}\bm{p}^{\text{tie}}_{i}\leq\bm{b}^{\text{tie}}_{i}\right.\right\}$ (29a)
^$\hat{\mathcal{E}}_{t}^{\text{bd}}:=\left\{\bm{x}_{t}^{\text{bd}}\left\lvert\,\left\|\left(\bm{E}_{t}^{\text{bd}}\right)^{-1}\left(\bm{x}_{t}^{\text{bd}}-\bm{e}_{t}^{\text{bd}}\right)\right\|\leq 1\right.\right\}$ (29b)
The whole dimension-decomposition-based flexibility aggregation algorithm is summarized in Algorithm 2.
Algorithm 2 Dimension-Decomposition-Based Flexibility Aggregation for a Regional Power Grid
1:  Analyze the temporal and spatial coupling relationships of boundary variables $\bm{x}^{\text{bd}}$. Categorize all the boundary variables into two types:
  • Temporal coupled variables $\bm{p}^{\text{tie}}_{i},\forall i\in\mathcal{I}^{\text{tie}}$
  • Single time slice coupled variables $\bm{x}^{\text{bd}}_{t},\forall t\in\mathcal{T}$
2:  Decompose the flexibility region into a series of lower-dimensional subspaces $\operatorname{space}(\bm{p}^{\text{tie}}_{i}),\forall i\in\mathcal{I}^{\text{tie}}$ and $\operatorname{space}(\bm{x}^{\text{bd}}_{t}),\forall t\in\mathcal{T}$. Apply shape templates to inner-approximate the projected regions in sub-spaces:
  • Use the polytopal template $\hat{\mathcal{G}}^{\text{tie}}_{i}$ (20) for $\operatorname{space}(\bm{p}^{\text{tie}}_{i})$
  • Use the ellipsoidal template ^$\hat{\mathcal{E}}_{t}^{\text{bd}}$ (21) for $\operatorname{space}(\bm{x}^{\text{bd}}_{t})$
Obtain the corresponding parameters $\tilde{\bm{b}}^{\text{tie}}_{i},\forall i\in\mathcal{I}^{\text{tie}}$ and $\tilde{\bm{E}}_{t}^{\text{bd}},\bm{e}_{t}^{\text{bd}},\forall t\in\mathcal{T}$, respectively.
3:  Reconstruct the inner-approximated flexibility region in the full space $\operatorname{space}(\bm{x}^{\text{bd}})$ by recombining the sub-regions $\hat{\mathcal{G}}^{\text{tie}}_{i},\forall i\in\mathcal{I}^{\text{tie}}$ and ^$\hat{\mathcal{E}}_{t}^{\text{bd}},\forall t\in\mathcal{T}$ according to (22). Obtain the corresponding parameters $\tilde{\bm{b}}^{\text{tie}}_{i},\forall i\in\mathcal{I}^{\text{tie}}$ and $\tilde{\bm{E}}_{t}^{\text{bd}},\bm{e}_{t}^{\text{bd}},\forall t\in\mathcal{T}$.
4:  Adjust the parameters $\tilde{\bm{b}}^{\text{tie}}_{i},\forall i\in\mathcal{I}^{\text{tie}}$ and $\tilde{\bm{E}}_{t}^{\text{bd}},\forall t\in\mathcal{T}$ using Algorithm 1 to ensure ^$\hat{\mathcal{R}}_{\text{L}}\subseteq\mathcal{R}_{\text{L}}$ is satisfied. Output the adjusted parameters as $\bm{b}^{\text{tie}}_{i},\forall i\in\mathcal{I}^{\text{tie}}$ and $\bm{E}_{t}^{\text{bd}},\forall t\in\mathcal{T}$.
5:  Return the final aggregated flexibility region for the regional power grid as defined in (28).